a bullet of mass 100g moving with a velocity of 240 m/s hits a target.After passing through the targetb10 cm thick ,it continued to move with a velocity of 160 m/s.Find the average resistance offered by the target.

Grade:9

1 Answers

Asiya Hassan
9 years ago
initial velocity(u)=240 m/s
Final velocity (v)=160m/s
Distance travelled=10cm=0.1m
v^2 – U^2 = 2as
a =v^2 -U^2/2s=(160)^2 – (240)^2/2*0.1= -16*10^4m/s^2
average resistance=Retardation offered by the target=16*10^4m/s^2